OpenStudy (sleepyjess):

A nick on the edge of a CD rotates to (−6, 5) during one song when represented graphically. What is the sine value of this function? 5 times square root 61 over 61 negative 6 times square root 61 over 61 negative square root 61 square root 5

OpenStudy (sleepyjess):

@jdoe0001

OpenStudy (jdoe0001):

\(\large {\begin{array}{llll} (-6&5)\\ (x&y)\\ (a&{\color{blue}{ b}}) \end{array}\qquad \begin{array}{llll} {\color{purple}{ c}}^2=a^2+b^2\implies {\color{purple}{ c}}=\sqrt{a^2+{\color{blue}{ b}}^2} \\ \quad \\ \bf sin(\theta)=\frac{{\color{blue}{ b}}}{{\color{purple}{ c}}} \end{array} }\)